# Knowledge Graph KGE Knowledge Graph Embedding

### 06 Mar 2026 v1

@@ 5 Experiments
Our experiments are designed to evaluate the performance gains provided by our caching system in both instance retrieval and concept learning tasks across multiple reasoners. To better understand the contribution of semantic awareness in our approach, we compare three configurations: (i) no cache, (ii) our semantic cache, and (iii) the non-semantic-aware variant of the cache (Non-semantic Cache). This cache follows the same storage and retrieval interface as our semantic cache, but does not analyze the syntactic or semantic structure of the concept expression. Unlike our semantic implementation, this baseline performs no decomposition into DL constructors, therefore corresponds to a naïve memoization mechanism and allows us to isolate the effect of semantic awareness from the effect of caching alone.
Since, to the best of our knowledge, no existing caching mechanisms have been specifically developed for class expression learning, nor even for instance retrieval, we use these three settings as our baselines to assess the impact of caching and, more importantly, the benefits of semantic-awareness in accelerating reasoning tasks.
The reasoners represent the state-of-the-art methods commonly used in the concept learning framework. In our evaluation, we selected the symbolic reasoners JFact [31], HermiT [11], Pellet [29], and its extended version Openllet [28]. Complementing these, we included an embedding-based reasoner (Ebr) [30], to explore a neural reasoning paradigm.
The evaluation involves two stages. Initially, we measure the performance of the reasoners without using our caching mechanism. Following this, we integrate our caching system into the experiments, varying the cache sizes to examine its impact. The cache size is relative to the total number of concepts generated for each dataset; for example, a cache size of $10\%$ implies that the cache can accommodate up to $10\%$ of the number of generated concepts on the corresponding dataset (see Table 2 to see the number of generated concepts on each dataset).
5.1 Experimental Setup
Concepts generation
Even though this algorithm can be easily extended to more expressive DL (e.g. $\mathcal{ALCIQ}$), we restrict our experiments to generate concepts of length up to 5 in the $\mathcal{ALC}$ DL (see Table 1 for syntax). The refinement operator described in [17] is used for this purpose. The total number of concepts generated for each dataset is summarized in Table 2.
@@ 5.1 Experimental Setup
Father 00006 004 01 03 000004 0090
We evaluate our caching system on five benchmark datasets: Father, Family, Carcinogenesis, Mutagenesis, and Vicodi [4]. The Father and Family datasets are relatively small (based on the number of individuals) and are commonly used for testing algorithms involving basic family and hierarchical relationships. To assess the scalability of our approach, we used larger and more complex datasets: Vicodi, Carcinogenesis, and Mutagenesis. These datasets are well-known in bioinformatics and cheminformatics, containing detailed information about chemical compounds and their biological effects, such as carcinogenicity and mutagenicity [16]. The key statistics of these datasets are presented in Table 2.
6 Results and Discussion
6.1 Application to Concept Retrieval
Runtime analysis
In Figure 3, the gray dotted line represents the baseline performance of each reasoner without caching. Since we reran the reasoner independently for each cache size setting, the gray line is not perfectly flat but shows slight variations. Across all reasoners, the integration of caching significantly reduces the run time, particularly at larger cache sizes. Notably, the LRU eviction strategy consistently delivers the best performance in minimizing run time. For example, when the cache has enough space to contain $80\%$ of the concepts, LRU enables dramatic run time reductions, outperforming other strategies. On the other hand, the performance of eviction strategies such as RP and MRU fluctuates, indicating less consistency across reasoners. LIFO also exhibits reasonable run time reductions but is outperformed by LRU in most cases.
Hit ratio analysis
Figure 3 illustrates the hit ratio performance on the Family dataset (all the other datasets showed similar hit ratio behaviour). As expected, larger cache sizes lead to higher hit ratios across all eviction strategies. LRU emerges as the most efficient strategy, although all reasoners achieve near-perfect hit ratios at higher cache allocations, LRU already shows high hit ratios at lower cache sizes (20%, 40%). In contrast, FIFO and MRU exhibit slower increases in hit ratios, suggesting less efficient utilization of cache space.
These results clearly indicate that our caching system enhances the performance of all reasoners, with LRU consistently outperforming other eviction strategies. This is evident in both reduced run times and increased hit ratios. This confirms the findings of O’neil et al. [23] who proved the superiority of LRU among different replacement policies. The scalability of our caching algorithm is demonstrated effectively on small datasets, setting the stage for its evaluation on larger datasets.
(a) Carcinogenesis
(b) Mutagenesis
(a) *
(b) *
(c) *
(d) *
(e) *
Figure 2: Run time performance vs. cache size for each reasoner on the Family dataset. The gray dotted line indicates the runtime of the reasoner without our cache.
(f) *
(g) *
(h) *
(i) *
(j) *
Figure 3: Hit score performance vs cache size for each reasoner on the Family dataset.
6.2 Application to Class Expression Learning
@@
To assess the effectiveness of our caching algorithm in CEL, we evaluated its impact on runtime performance across four datasets and several state-of-the-art CEL algorithms. Specifically, we included three top-down refinement-based learners—OCEL, CELOE, and CLIP—as well as EvoLearner, which follows a different strategy (see Section 2 for details). All learning problems and corresponding sets of positive and negative examples were generated following the protocol introduced by [8], and are available in the supplementary material. In this experiment, the maximum storage capacity of the cache has been set to 1024 concepts.
The results presented in Figure 4 demonstrate the substantial impact our cache algorithm has on accelerating the concept learning process, particularly for top-down refinement-based learners such as OCEL, CELOE, and CLIP. Among these, OCEL benefits the most. For instance, on the Carcinogenesis dataset, the average runtime of OCEL drops dramatically from over 100 seconds without the cache to under 20 seconds with it. Similarly, on the Mutagenesis dataset, the runtime decreases from more than 25 seconds to approximately 12 seconds. These improvements are consistent with OCEL’s reliance on expensive instance retrieval operations, which our cache significantly optimizes by avoiding redundant reasoning through semantic reuse.
By contrast, the non-semantic cache shows little to no improvement for most learners and, in some cases, even leads to worse performance. This degradation stems from the cache being quickly filled with syntactically distinct but semantically irrelevant concepts, overwhelming the memory with uninformative entries. As a result, many cache lookups fail or return unhelpful data, causing additional overhead rather than reducing it. These findings highlight the importance of semantic awareness: without it, caching can hinder rather than help.
CELOE and CLIP also show measurable runtime reductions across datasets, albeit to a lesser extent. On the Vicodi dataset, CELOE’s average runtime drops from roughly 135 seconds to about 100 seconds. These learners, while still based on top-down refinement strategies, tend to produce fewer costly retrievals or shallower refinement chains, which may explain why the performance gains, though present, are not as dramatic as with OCEL. Nevertheless, the consistent improvements observed with CELOE and CLIP confirm that our caching approach generalizes well across multiple concept-learning systems that employ refinement operators.
However, EvoLearner shows no benefit from our caching mechanism. This result aligns with EvoLearner’s nature, which does not follow a top-down refinement strategy. Instead, EvoLearner begins the learning process by sampling a large population of candidate concepts and their associated individuals upfront. This sampling-based design inherently reduces its reliance on repeated instance retrieval during the search process. Furthermore, the large initial population leads to a wide set of unique candidate concepts whose instance sets must be computed and stored early on. Attempting to cache all this information is not only memory-intensive but also largely inefficient, as the likelihood of reusing these exact expressions is low in an evolutionary setting. As a result, our cache provides limited benefit in EvoLearner’s workflow, explaining the difference in runtime between the cached and non-cached variants.

### 12 Mar 2026 v2

@@ 5 Experiments
Tasks and Baselines
Our experiments evaluate the impact of our caching system on concept retrieval and concept learning tasks. We compare three configurations: (i) no cache, (ii) our semantic cache, and (iii) a non-semantic cache. The latter uses the same storage interface but does not exploit the syntactic or semantic structure of concept expressions, effectively corresponding to naïve memoization. To the best of our knowledge, no caching mechanism has been specifically proposed for class expression learning or concept retrieval. We therefore use these three configurations as baselines to assess both the benefits of caching and the additional gains from semantic reasoning.
For concept retrieval, concepts with a maximum length of 10 are generated by randomly sampling concepts from the refinement tree produced by any concept learner. The total number of generated concepts per dataset is summarized in Table 2. For concept learning, we rely on existing learning problems (LPs), which are provided in the supplementary material.
We evaluate our approach using several state-of-the-art reasoners commonly used in concept retrieval frameworks. The symbolic reasoners include JFact [31], HermiT [11], Pellet [29], and its extended version Openllet [28]. In addition, we include a neural reasoning approach (Ebr) [30] to investigate the effectiveness of the cache in a neuro-symbolic setting.
Our evaluation proceeds in two stages. First, we measure the performance of the reasoners without caching to establish a baseline. We then enable the cache and vary its size to analyze its impact on runtime and retrieval efficiency. The cache size is defined relative to the number of concepts generated for a dataset. For instance, a cache size of $10\%$ indicates that the cache can store up to $10\%$ of the total number of generated concepts (see Table 2). In addition to runtime performance, we also measure the physical memory consumption of the cache for both the semantic and non-semantic variants in order to quantify the average memory requirements of the caching mechanism.
We also analyze the impact of cache initialization. On the small Family dataset, we evaluate the cache both with and without the initialization phase to measure its contribution. For larger datasets, however, the cache is used only with initialization enabled, as this configuration consistently yields better performance and more stable behavior in practice.
Hardware
All experiments were conducted on a virtual machine equipped with 2 NVIDIA H100-80C GPUs (80GB each) and 32 Intel Xeon Platinum CPUs with 31GB of memory. The GPUs are required for batch embedding computation in the neuro-symbolic reasoner Ebr.
@@
We evaluate our cache on four benchmark datasets: Family, Carcinogenesis, Mutagenesis, and Vicodi [4]. Family is relatively small and commonly used for testing algorithms involving family relationships. To evaluate scalability, we also use larger datasets: Vicodi, Carcinogenesis, and Mutagenesis, which are widely used in bioinformatics and cheminformatics and contain detailed information about chemical compounds and their biological properties [16]. Key dataset statistics are reported in Table 2.
@@
6 Results and Discussion
6.1 Results on Concept Retrieval
Runtime analysis
Figures 4 and 6 present the runtime results for concept retrieval on the Family dataset with and without cache initialization, respectively. The gray dotted line represents the baseline runtime of each reasoner without caching. Since the reasoner is executed independently for each cache size configuration, the baseline line is not perfectly flat and exhibits minor variations.
Across all reasoners, integrating the cache significantly reduces runtime, particularly as the cache size increases. Among the evaluated eviction strategies, LRU consistently provides the best performance. For instance, when the cache can store up to $80\%$ of the generated concepts, LRU yields the largest runtime reductions across all reasoners. In contrast, strategies such as RP and MRU show higher variability and less consistent improvements. LIFO also reduces runtime but generally remains slightly less effective than LRU.
Comparing warm and cold cache settings, the warm cache (initialized) leads to a more stable and faster decrease in runtime as the cache size grows. This improvement comes at the cost of a slightly higher runtime at the beginning of the execution due to the initialization phase. Overall, once the cache becomes populated, the warm configuration provides more consistent performance gains.
Hit ratio analysis
Figures 4 and 6 show the hit ratio results on the Family dataset for the warm and cold cache settings, respectively. The other datasets exhibit similar trends.
As cache size increases, the hit ratio improves for all eviction strategies. LRU again demonstrates the most efficient behavior, achieving high hit ratios even at relatively small cache sizes (e.g., $20\%$ and $40\%$). In contrast, FIFO and MRU display slower improvements, indicating less efficient use of the available cache space. While the final hit ratios of warm and cold caches are similar at large cache sizes, the warm cache generally reaches high hit ratios faster and exhibits more stable behavior.
Overall, these results demonstrate that the proposed caching system substantially improves the performance of all evaluated reasoners. LRU consistently provides the best trade-off between runtime reduction and cache efficiency in both warm and cold cache settings. This observation aligns with the theoretical results of O’Neil et al. [23], which establish the optimality of LRU among common replacement policies. The results on the Family dataset, therefore, validate the effectiveness of our caching approach and motivate its evaluation on larger and more complex datasets.
(a) Carcinogenesis
(b) Mutagenesis
(a) *
(b) *
(c) *
(d) *
Figure 3: Run time performance vs. cache size for each reasoner on the Family dataset with cache initialization. The gray dotted line indicates the runtime of the reasoner without our cache.
(e) *
(f) *
(g) *
(h) *
Figure 4: Hit score performance vs cache size for each reasoner on the Family dataset with cache initialization. Higher values indicate more cache hits during retrieval.
(a) *
(b) *
(c) *
(d) *
Figure 5: Run time performance vs. cache size for each reasoner on the Family dataset without cache initialization. The gray dotted line indicates the runtime of the reasoner without our cache.
(e) *
(f) *
(g) *
(h) *
Figure 6: Hit score performance vs cache size for each reasoner on the Family dataset without cache initialization. Higher values indicate more cache hits during retrieval.
Carcinogenesis
(a) *
Vicodi
(b) *
Mutagenesis
(c) *
Family
(d) *
Figure 7: Average physical memory usage of concept learners using the semantic cache, and the non-semantic cache baseline across different datasets.
6.2 Results on Class Expression Learning
@@
To evaluate the impact of our caching algorithm on CEL, we measured runtime performance across four datasets using several state-of-the-art learners. We considered three top-down refinement-based approaches—OCEL, CELOE, and CLIP—as well as EvoLearner, which follows an evolutionary strategy (see Section 2). Learning problems and corresponding positive and negative examples were generated following the protocol of [8]. In all experiments, the maximum cache capacity was set to 1024 concepts.
The results in Figure 1 show that our semantic cache substantially accelerates concept learning for refinement-based learners. The most pronounced improvements are observed for OCEL. For example, on the Carcinogenesis dataset, the average runtime decreases from more than 100 seconds without caching to under 20 seconds with the semantic cache. Similar reductions are observed on Mutagenesis and Vicodi. These gains are expected since refinement-based learners repeatedly perform concept retrieval operations, which our cache accelerates by reusing previously computed results.
CELOE and CLIP also benefit from the cache, although to a lesser extent. For instance, on Vicodi, the runtime of CELOE decreases from roughly 135 seconds to around 100 seconds. While these learners still rely on refinement operators, they typically generate fewer expensive retrieval queries, which explains the smaller but consistent improvements.
In contrast, the non-semantic cache provides little or no improvement and occasionally degrades performance. Because it stores concepts without considering their semantic structure, the cache quickly fills with syntactically distinct expressions that provide little opportunity for reuse. As a result, cache lookups rarely produce useful results and introduce additional overhead. This behavior highlights the importance of semantic awareness in enabling effective reuse.
EvoLearner does not benefit from the cache. Unlike refinement-based learners, it generates a large population of candidate concepts at the beginning of the search and evaluates them upfront. Consequently, concept retrieval is performed mostly once per concept, leaving little opportunity for reuse through caching.
Figure 7 reports the average physical memory consumption of the semantic and non-semantic caches across datasets. The results show that both variants exhibit similar memory usage, indicating that the semantic reasoning layer does not introduce significant additional overhead. Memory consumption mainly depends on the number of stored concepts and the size of their instance sets rather than on the semantic processing itself.
@@ 7 Conclusion
Acknowledgements
This work was funded by the EU’s Horizon Europe programme (Marie Skłodowska-Curie grant No. 101073307) and supported by the Ministry of Culture and Science of North Rhine-Westphalia (MKW NRW) through the WHALE project (LFN 1-04) and the SAIL project (grant No. NW21-059D).